Abstract

Traces and Invariants for Non-Compact Manifolds

We define a general trace for complete Riemannian manifolds of bounded geometry. This trace takes values in the Haefliger functions of the manifold and its existence allows the extension of the classical theory of invariants for compact manifolds to complete Riemannian manifolds of bounded geometry.